# Poisson Hail on a Wireless Ground

### v1

@@ Abstract
This paper defines a new model which incorporates three key ingredients of a large class of wireless communication systems: (1) spatial interactions through interference, (2) dynamics of the queueing type, with users joining and leaving, and (3) carrier sensing and collision avoidance as used in, e.g., WiFi. In systems using (3), rather than directly accessing the shared resources upon arrival, a customer is considerate and waits to access them until nearby users in service have left. This new model can be seen as a missing piece of a larger puzzle that contains such dynamics as spatial birth-and-death processes, the Poisson-Hail model, and wireless dynamics as key other pieces. It is shown that, under natural assumptions, this model can be represented as a Markov process on the space of counting measures. The main results are then two-fold. The first is on the shape of the stability region and, more precisely, on the characterization of the critical value of the arrival rate that separates stability from instability. The second is of a more qualitative or perhaps even ethical nature. There is evidence that for natural values of the system parameters, the implementation of sensing and collision avoidance stabilizes a system that would be unstable if immediate access to the shared resources would be granted. In other words, for these parameters, renouncing greedy access makes sharing sustainable, whereas indulging in greedy access kills the system.
1 Taxonomy and Motivations
@@
We discuss here four basic classes of spatial queuing dynamics. In each case, there is a time-space Poisson process of customer arrivals on a compact ${\cal X}\subset\mathbb{R}^{d}$. Arrivals are characterized by a time $t\in\mathbb{R}$ and a locus $x\in\cal X$. Each arrival also comes with a service demand (or file size) $h\in{\mathbb{R}}^{+}$, that may vary from customer to customer. The compact $\cal X$ can be seen as the server. The intensity measure of this Poisson process is assumed to be of the form $\lambda\mathrm{d}t\mu(\mathrm{d}x)$, where $\mu$ is a measure on $\mathcal{X}$ with $\mu(\cal X)<\infty$.
@@
It is best to start with the Poisson Hail on a Hot Ground (PHHG) dynamics introduced in [1]. In this model, in addition to their service demand, each arrival at locus $x$ comes with an exclusion set $S\subset\mathbb{R}^{d}$. A typical instance, which is used in this introduction, is that where $S=B(x,R)$, namely the closed ball with radius $R$ centered at $x$, where $R>0$ is a fixed constant or a random variable. Two arrivals are said to be conflicting if their exclusion sets intersect. No two conflicting customers can be served simultaneously. The service of a customer with arrival time $t$ starts when all conflicting customers that arrived before $t$ have left the system. This can be seen as a local First-Come-First-Serve (FCFS) rule. The service speed is constant. Here is what explains the terminology. The ball of a customer is the footprint of the hailstone associated with this customer and its service demand is the height of this hailstone. The temperature of the compact $\cal X$ (subset of the ground) is constant so that the hailstones touching the ground melt at constant speed, whereas the others form a (cold) heap-shaped spatial queue.
@@
The second dynamics is that obtained from the PHHG when setting $S=\{x\}$ (or $R=0$), namely a hailstone arriving at $x$ is an ice-stick at $x$. In this case, when the space marginal measure $\mu$ is absolutely continuous w.r.t. Lebesgue measure, no two customers ever conflict a.s., so that each stick receives immediate attention and melts at constant speed. This model is the spatial birth and death (SBD) process introduced in [2]. If the space marginal is atomic, we get a queuing network made of a finite collection of independent M/M/1 queues located at the atoms of $\mu$.
@@
The third model is the wireless spatial birth and death (WSBD) process which is a variant of the SBD introduced in [3]. The only difference with the SBD is in the service rate (or melting speed) of sticks. Rather than constant speed, the service rate of a customer at $x$ is proportional to the Shannon rate at $x$ when considering all other concurrent customers as interferers and treating interference as noise. The rationale for this service rate comes from information theory and wireless communications [4]. The service discipline is hence immediate attention but with varying service rates across time and space.
@@
The present paper is focused on a model which is new to the best of our knowledge and that we will refer to as the Poisson hail on a wireless ground (PHWG) model. It is meant to study arrivals with (non-trivial) exclusion sets and varying service rates across time and space. In the wireless setting, the exclusion set serves as a protection zone around wireless transmissions, which is a way to control the medium access similar to what is introduced in carrier-sense multiple access protocols (see below). The dynamics can be described as follows:
@@
  • Shannon rate: The service rate of each active hailstone at any time $t$ is its Shannon rate, which is a deterministic function of the configuration of all active hailstones at time $t$ (see Eq (1)). Note that at a given time $t$, this rate may be different for two customers being served at that time.
@@
  • The variant of the PHHG model when the temperature of the ground varies with time and space according to Shannon’s capacity formula, namely the melting speed at time $t$ of a hailstone touching the ground and centered at $x$ is given by the Shannon rate, which depends on all hailstones touching the ground at that time.
@@
  • The variant of the WSBD model where, rather than starting transmitting immediately upon arrival, an arrival is required to apply certain priority rules based on their exclusion sets and which are those of the PHHG model.
@@
The rationale for introducing this model goes beyond the sake of completeness. As suggested by (parts of) the terminology, the practical motivations are to be found in the analysis of wireless communications (see Section 2) at the system level. System-level properties include, for instance, the capacity region of the queuing system of the mean system time of a typical customer. By capacity, we mean the queuing theoretic capacity, namely the set of values for $\lambda$ such that the dynamics in question is stable (say positive recurrent under a proper Markov representation) when fixing all other system parameters. The fact that this stability region of PHWG is of the form $\lambda<\lambda_{c}$ when all other parameters are fixed is one of the main results established in the present paper, together with a constructive representation of $\lambda_{c}$. As we shall see, there are system parameter regions for which the WSBD is unstable, whereas the PHWG is stable. In more concrete terms, in the stability region in question, the queuing delays incurred by some customers due to the exclusion/protection rule in fact allow one to get finite system times in the steady state, whereas stationary system times are infinite in the situation with immediate access to the channel ($R=0$). We find it interesting that refraining from immediate access to a system could stabilize the system in question. This can be seen as a strong argument in favor of protocols of the CSMA-CA-type (see Section 2), which control the access to the shared medium, in order to prevent that nearby customers access that shared channel at the same time. In the stable case, we also analyze the stationary distribution of the waiting or system time of a typical customer and its dependence on $R$. Even if wireless motivations are essential, we will show (see Section 6) that this class of models has, in fact, a much broader scope and potential set of applications.
2 Contributions
@@
The SBD dynamics was first studied in [2] and was then the object of numerous further studies. The PHHG dynamics was introduced in [1]. The stability condition is known. The steady state can be constructed for both compact and infinite spaces. Both dynamics are rather well understood. The WSBD model was introduced in [3] for the compact and continuous space case. The infinite discrete space case was analyzed in [5], where the stability conditions were obtained and a closed form for the steady state queue size was derived. The infinite continuous state space case is unsolved at this stage, to the best of our knowledge.
@@
As explained above, the PHWG model, which is the central topic of the present paper, admits both WSBD and PHHG (and hence SBS) as special cases. In view of the state of the art for these special cases, we will restrict ourselves to the compact state space case. The first objectives are the construction of the steady state and the identification of the stability region of PHWG. Our main result on the matter is the proof of the fact that, when fixing all parameters of the model except $\lambda$, there exists a critical value $\lambda_{c}$ such that if $\lambda<\lambda_{c}$, the PHWG dynamics is a positive recurrent continuous Markov chain on the space of counting measures, whereas it is transient if $\lambda>\lambda_{c}$. Another aim is the comparison of PHWG with all the dynamics listed above. One specific example is the identification of the parameter region where (in some sense to be defined more precisely below) PHWG does better than WSBD. Since WSBD is obtained when setting $R=0$ in PHWG, a simple instance is the identification of the set of parameters other than $R$ such that there is an optimum of $\lambda_{c}$ for $R=R^{*}$ with $R^{*}\neq 0$, and the determination of $R^{*}$. A third aim of the paper is the extension of this model to spatial birth and death with interaction (or death) functions other than that associated with the Shannon rate. As already mentioned, we show that there is a wide class of dynamics where the approach and results developed here extend in a natural way.
2.1 Mathematical Approach
@@
A key step in the construction of the steady state and the identification of the stability region is the identification of arrival blocks that form i.i.d sequences. A sufficient condition to construct such blocks is the existence of what we call zigzag events in the exclusion sets. The simplest such event consists in a single exclusion set covering the whole space. As we show below, more refined zigzag events occur whenever each exclusion set contains a ball of deterministic radius $R$ where $R$ is any positive constant common to all arrivals. More refined events can be considered, linked to the rank of tropical matrices.
2.2 Implications for Wireless Networks
@@
The main novelty of the paper in terms of wireless communications is in the analysis of the dynamics of the spatial properties of carrier-sense multiple access with collision avoidance (CSMA/CA). The principle of CSMA/CA consists in sensing the shared medium and in refraining accessing it if there is a nearby transmitter already using it. There is an immense literature on this multiple access scheme. Its spatial but static analysis was first considered using a Matérn point process heuristic in [6], without analysis of dynamic stability. [7] defines and analyzes the conditions for $\epsilon$-stability, where at most $\epsilon$-fraction of wireless queues are instable. Mean-field heuristics of the dynamics were proposed [8], with an approach ignoring the spatial nature of the problem and in [8] when taking this feature into account. To the best of our knowledge, the present paper is the first to propose an exact (no mean-field approximation) analysis of the spatial properties of CSMA/CA type dynamics.
@@
Let us stress that the version of the collision avoidance mechanism studied in the present paper is quite different from that used in the current versions of the CSMA-CA protocol. The main common feature is the implementation of an exclusion principle which forbids that two (or more customers) with a mutual distance smaller than a given value could transmit simultaneously. The FCSF mechanism used in the paper in order to implement this is quite different from what is currently done.
2.3 Paper Organization
@@
The paper is structured as follows: Section 3 defines the dynamics. Section 4 establishes the main stability result. Section 5 gives simulation and comparison results. Section 6 gives the class of Spatial Interaction Birth and Death (SIBD) processes to which our main results can be extended. Section 7 contains various discussions and a list of open problems. Finally, the appendix gathers the proofs of our two main theorems.
3 Description of the Dynamics
3.1 Arrival Process
We model the arrival of customers as a stationary marked Poisson space-time process of intensity $\lambda$ on $\cal X\times\mathbb{R}$, denoted by $\mathcal{A}=\{\left((x_{i},t_{i}),(S_{i},h_{i})\right)\}_{i\in\mathbb{N}^{+}}$, where $(x_{i},t_{i})$ denotes the location and arrival time of the $i$-th customer, ordered by increasing time of arrival. Each arrival has a mark (or “hailstone”) $(S_{i},h_{i})$, where $S_{i}\subset{\cal X}$ is the “footprint” (e.g. a ball) and $h_{i}\in\mathbb{R}^{+}$ is the “height” of the hailstone. The random sequence of marks $\{(S_{i},h_{i})\}$ is assumed to be i.i.d. and not to depend on $\{(x_{i},t_{i})\}$. $S_{i}$ and $h_{i}$ need not be independent. In what follows, we omit the index $i$ if it is not needed. For instance, $S$ is a random variable with the same distribution as $S_{1}$. We assume that $h$ has a finite mean. In the wireless setting, the $i$-th customer is a wireless link where the transmitter and receiver are co-located. This simplification, which makes sense when the distance from transmitter to receiver is small compared to the distances between links, is that used in [3].
The dynamics is depicted in Figure 1. A discrete-time version is illustrated for simplicity.

### v2

@@ Abstract
This paper defines a new model which incorporates three key ingredients of a large class of wireless communication systems: (1) spatial interactions through interference, (2) dynamics of the queueing type, with users joining and leaving, and (3) carrier sensing and collision avoidance as used in, e.g., WiFi. In systems using (3), rather than directly accessing the shared resources upon arrival, a customer is considerate and waits to access them until nearby users in service have left. This new model can be seen as a missing piece of a larger puzzle that contains such dynamics as spatial birth-and-death processes, the Poisson-Hail model, and wireless dynamics as key other pieces. It is shown that, under natural assumptions, this model can be represented as a Markov process on the space of counting measures. The main results are then two-fold. The first result is on the shape of the stability region and, more precisely, on the characterization of the critical value of the arrival rate that separates stability from instability. The second result is of a more qualitative or perhaps even ethical nature. There is evidence that for natural values of the system parameters, the implementation of sensing and the delayed access for collision avoidance can stabilize a system that would be unstable if immediate access to the shared resources would be granted. In other words, for these parameters, renouncing greedy access makes sharing sustainable, whereas indulging in greedy access kills the system.
I Introduction
I-A Background
Spatial regulation is a central issue in the access to the wireless network medium. It is somewhat surprising that such a regulation is fundamental in carrier-sense multiple access with collision avoidance (CSMA-CA) protocols, for instance, where close-by simultaneous access (“collision”) to the shared medium is prevented (“avoided”), whereas it is only present through optional spatial reuse mechanisms in most 3GPP medium access schemes. The rationale for such a regulation is, of course, the mitigation of interference: when treating interference as noise, a wireless system sees its performance significantly degrade when the power of interference is high. The general question is that of the optimal trade-off between the gains obtained on the communication rates of scheduled users by delaying medium access of the other users, and the negative consequences of the delays incurred by the latter.
This question has been central since the 70’s. One of the difficulties is the definition of metrics that incorporate in a rigorous way the spatial components that are inherent in this class of problems and the dynamical system components of the problem. Among the latter, we would quote the epochs at which packets to be transmitted by their wireless links are generated by users, the delays which are incurred by these packets due to spatial regulation and more generally due to local queuing, as well of course as delays which are incurred by them for transmission on the wireless links once they are scheduled.
The present paper proposes a new model that is meant to address this old question in a new way. This model integrates spatial arrival processes, spatial regulation as described above, as well as an exact representation of scheduled wireless link transmission rates based on classical information-theoretic paradigms. In contrast to earlier approaches, this new model gives an exact representation of the resulting dynamics and is amenable to a mathematical analysis that is not based on an approximation (there is no mean-field approximation in particular). It allows us to define a rigorous metric, which is the dynamic stability capacity of the system. This new model is used to prove that appropriate spatial regulation may allow one to increase this dynamic stability capacity. This is the central result of the paper, together with the definition of the setting allowing one to formalize the problem.
I-B Taxonomy and Motivations
@@
We discuss here four basic classes of spatial queuing dynamics. In each case, there is a space-time Poisson process of customer arrivals on a compact ${\cal X}\subset\mathbb{R}^{d}$. Arrivals are characterized by a time $t\in\mathbb{R}$ and a locus $x\in\cal X$. Each arrival also comes with a service demand (or file size) $h\in{\mathbb{R}}^{+}$, that may vary from customer to customer. The compact $\cal X$ can be seen as the server. The intensity measure of this Poisson process is assumed to be of the form $\lambda\mathrm{d}tm(\mathrm{d}x)$, where $\lambda>0$ and $m$ is a measure on $\mathcal{X}$ with $m(\cal X)<\infty$.
@@
It is best to start with the Poisson Hail on a Hot Ground (PHHG) dynamics introduced in [2]. In this model, in addition to their service demand, each arrival at locus $x$ comes with an exclusion set $S\subset\mathbb{R}^{d}$. A typical instance, which is used in this introduction, is that where $S=B(x,R)$, namely the closed ball with radius $R$ centered at $x$, where $R>0$ is a fixed constant or a random variable. Two arrivals are said to be conflicting if their exclusion sets intersect. No two conflicting customers can be served simultaneously. The service of a customer with arrival time $t$ starts when all conflicting customers that arrived before $t$ have left the system. This can be seen as a local First-Come-First-Served (FCFS) rule. The service speed is constant. Here is what explains the terminology. The ball of a customer can be seen as the footprint of the hailstone associated with this customer and its service demand is the height of this hailstone. The temperature of the compact $\cal X$ (subset of the ground) is constant so that the hailstones that touch the ground melt at constant speed, whereas the others form a (cold) pile-shaped spatial queue.
@@
The second dynamics is that obtained from the PHHG when setting $S=\{x\}$ (or $R=0$), namely a hailstone arriving at $x$ is an ice-stick at $x$. In this case, when the measure $m$ is absolutely continuous w.r.t. Lebesgue measure, no two customers ever conflict a.s., so that each stick receives immediate attention and melts at constant speed. This model is the spatial birth-and-death (SBD) process introduced by Preston in [15]. If the space marginal is atomic, we get a queuing network consisting of a finite collection of independent FCFS M/M/1 queues located at the atoms of $m$ if the height variables are exponential and i.i.d.
@@
The third model is the wireless spatial birth and death (WSBD) process which is a variant of the SBD introduced in [17]. The only difference from the SBD is in the service rate (or melting speed) of the sticks. Rather than a constant speed, the service rate of a customer at $x$ is proportional to the Shannon rate at $x$ when considering all other concurrent customers as interferers and treating interference as noise. The rationale for this service rate comes from information theory and wireless communications [18]. The service discipline is hence with immediate attention but with varying service rates across time and space.
@@
The present paper is focused on a model that is new, to the best of our knowledge, and that we will refer to as the Poisson hail on a Wireless Ground (PHWG) model. It is meant to study arrivals with (non-trivial) exclusion sets and varying service rates across time and space. In the wireless setting, the exclusion set serves as a protection zone around wireless transmissions, which is a way to control the medium access similar to what is introduced in carrier-sense multiple access protocols (see below). The dynamics can be described in more detail as follows:
@@
  • Shannon rate: The service rate of each active hailstone at any time $t$ is its Shannon rate, which is a deterministic function of the configuration of all active hailstones at time $t$ (see Eq (2)). Note that, at a given time $t$, this rate may be different for two customers being served at that time.
@@
  • A variant of the PHHG model when the temperature of the ground varies with time and space according to Shannon’s capacity formula, namely the melting speed at time $t$ of a hailstone touching the ground and centered at $x$ is given by the Shannon rate, which depends on the locations of all hailstones touching the ground at that time.
@@
  • A variant of the WSBD model where, rather than starting transmitting immediately upon arrival, an arrival is required to apply certain priority rules based on their exclusion sets and which are those of the PHHG model.
@@
The rationale for introducing this model goes beyond the sake of completeness. As suggested by (parts of) the terminology, our practical motivations are to be found in the analysis of wireless communications (see Section I-C) at the system level. System-level properties include, for instance, the capacity region of the queuing system or the mean system time of the typical customer. By capacity, we mean the queuing theoretic capacity, namely the set of values for $\lambda$ such that the dynamics in question is stable (say positive recurrent under a proper Markov representation), when fixing all other system parameters. The fact that this stability region of PHWG is of the form $\lambda<\lambda_{c}$ when all other parameters are fixed is one of the main results established in the present paper, together with a constructive representation of $\lambda_{c}$. As we shall see, there are system parameter regions for which the WSBD is unstable, whereas the PHWG is stable. In more concrete terms, in the stability region in question, the queuing delays incurred by some customers due to the exclusion/protection rule, in fact, allow one to get finite system times in the steady state, whereas stationary system times are infinite in the situation with immediate access to the channel ($R=0$). We find it interesting that refraining from immediate access to a system could stabilize the system in question. This can be seen as a strong argument in favor of protocols of the CSMA-CA-type (see Section I-C), which control the access to the shared medium, in order to prevent nearby customers from accessing that shared channel at the same time. In the stable case, we also discuss the stationary distribution of the waiting or system time of the typical customer and its dependence on $R$. Even if wireless motivations are essential, we will show (see Section V) that this class of models has, in fact, a much broader scope and potential set of applications.
I-C Contributions
@@
The SBD dynamics was first studied in [15] and was then the subject of numerous further studies. The PHHG dynamics was introduced in [2]. The stability condition is known. The steady state can be constructed for both compact and infinite spaces. Both dynamics are well understood. The WSBD model was introduced in [17] for the compact and continuous space case. A model with the infinite discrete space case was analyzed in [16], where the stability conditions were obtained and a closed form was derived for the steady state queue size. The infinite continuous state space case is unsolved at this stage, to the best of our knowledge.
@@
As explained above, the PHWG model, which is the central topic of the present paper, admits both WSBD and PHHG (and hence SBD) as special cases. In view of the state of the art for these special cases, we will restrict ourselves to the compact state space case. The first contributions of the paper are the construction of the steady state and the identification of the stability region of PHWG. Our main result on this matter is presented in Theorems 1 and 2: we prove that, when fixing all parameters of the model except $\lambda$, there exists a critical value $\lambda_{c}$ such that if $\lambda<\lambda_{c}$, the PHWG dynamics is a positive recurrent continuous Markov chain in the space of counting measures, while it is transient if $\lambda>\lambda_{c}$. Another contribution is the comparison of PHWG with the dynamics listed above. We study two examples in Section IV, one in the continuous-space setting and one in the discrete-space setting, to show that PHWG outperforms WSBD for natural values of the system parameters. Since WSBD is obtained when setting $R=0$ in PHWG, a simple instance is the identification of an optimum of $\lambda_{c}$ for $R=R^{*}$ with $R^{*}\neq 0$, and the determination of $R^{*}$ through numerical evaluations. A third outcome of the paper is the identification of extensions of this model to spatial birth and death with interaction (or death) functions other than that associated with the Shannon rate. We show that there is a wide class of dynamics where the approach and results developed here extend in a natural way.
I-C1 Mathematical Approach
@@
A key step in the construction of the steady state and the identification of the stability region is the identification of arrival blocks that form i.i.d. sequences. A sufficient condition to construct such blocks is the existence of what we call zigzag events. The simplest such event consists of a single exclusion set that covers the entire space. As we show below, more refined zigzag events occur whenever each exclusion set contains a ball of deterministic radius $R$ where $R$ is any positive constant common to all arrivals. More refined events can be considered, linked to the rank of tropical matrices.
I-C2 Implications for Wireless Networks
@@
The main novelty of the paper in terms of wireless communications is in the analysis of the dynamics of the spatial properties of carrier-sense multiple access with collision avoidance (CSMA/CA). The principle of CSMA/CA consists in sensing the shared medium and refraining from accessing it if there is a nearby transmitter already using it. There is a great deal of literature on this multiple access scheme. Its spatial but static analysis was first considered using a Matérn point process heuristic in [14], without analysis of dynamic stability. [7] defines and analyzes the conditions for $\epsilon$-stability, where at most $\epsilon$-fraction of wireless queues are instable. Mean-field heuristics associated with the dynamics were proposed [13], with an approach ignoring the spatial nature of the problem and in [6] when taking this feature into account. To the best of our knowledge, the present paper is the first to propose an exact (no mean-field approximation) analysis of the spatial properties of CSMA/CA type dynamics.
@@
It is important to stress that the version of the collision avoidance mechanism studied in the present paper is quite different from that used in the current versions of the CSMA-CA protocol. The main common feature is the implementation of an exclusion principle which forbids that two (or more) customers with a mutual distance smaller than a given value could transmit simultaneously. The local FCFS mechanism used in the paper in order to implement this is quite different from what is currently done.
I-D Paper Organization
@@
The paper is structured as follows: Section II defines the dynamics. Section III establishes the main stability result. Section IV gathers simulation and comparison results. Section V gives the class of Spatial Interaction Birth and Death (SIBD) processes to which our main results can be extended. Section VI contains various discussions and a list of open problems. Finally, the appendix gathers the proofs of our two main theorems.
II Description of the Dynamics
The aim of this section is to show that the sample path of the random dynamics described informally in the last section can be constructed by induction. The main outcome of the construction is that there are no accumulation points of jumps in the dynamics so that the latter is well defined for all positive times for any given initial condition.
II-A Arrival Process
The arrival process is best described as a stationary marked Poisson point process on $\mathbb{R}$ with i.i.d. marks on $\mathcal{X}\times\mathbb{R}^{2}$. The intensity of the arrival process is denoted by $\Lambda>0$. The probability distribution of the marks is denoted by $\mu\times\nu$, where $\mu$ is a probability measure on $\mathcal{X}$ and $\nu$ is a probability distribution on $\mathbb{R}^{2}$. This marked point process is denoted by $\mathcal{A}=\{\left((t_{i}),(x_{i},h_{i},R_{i})\right)\}_{i\in\mathbb{N}^{+}}$, where $t_{i}$ denotes the arrival time of the $i$-th customer (arrivals are ordered by increasing arrival times), $x_{i}$ the arrival locus of the $i$-th customer, $h_{i}$ the service time of the $i$-th customer, and $R_{i}$ its exclusion radius. The interpretation is that the $i$-th arrival comes as a “hailstone” $(S_{i},h_{i})$, where $S_{i}=B(x_{i},R_{i})$ is the “footprint” (which is here closed ball of radius) and $h_{i}\in\mathbb{R}^{+}$ is the “height” of the hailstone. The random sequence of pairs $\{(h_{i},R_{i})\}$ is assumed to be i.i.d. and not depend on $\{(x_{i},t_{i})\}$. However, $R_{i}$ and $h_{i}$ need not be independent. They are assumed to have a joint distribution $\nu$ on $\mathbb{R}^{2}$. This formulation is equivalent to the informal models of the introduction with
$\Lambda=\lambda m(\mathcal{X})\quad\mbox{and}\quad\mu(\cdot)=\frac{m(\cdot)}{m(\mathcal{X})}.$ (1)
In what follows, we will often omit the index $i$ if it is not needed. More precisely, $(x,h,R)$ will denote an independent copy of $(x_{1},h_{1},R_{1})$ and $S$ will denote the set $S=B(x,R)$. In the wireless setting, the $i$-th customer is a wireless link where the transmitter and receiver are co-located. This simplification makes sense when the distance from transmitter to receiver is small compared to the distances between links.
The dynamics is depicted in Figure 1.